Light strikes your cornea with an angle of incidence of 35.0°. The index of refraction for your cornea is 1.376, for air 1.000, for water 1.333.

What is the angle of refraction of this light when the eye is in air?

In water?
Answer #1

Given :-

n1 = 1

n2 = 1.376

n3 = 1.333

theta1 = 35 degree

from snell's law

n1*sin(theta)1 = n2*sin(theta)2

sin(theta)2 = (1 x sin(35)) / 1.376 = 0.4168433

(theta)2 = 24.6 degree .........ans(1) for the eye

n1*sin(theta)1 = n3*sin(theta)2

sin(theta)2 = (1 x sin(35)) / 1.333 = 0.4168433

(theta)2 = 25.4 degree .........ans(2) for the water